# Towards Trustworthy Knowledge Graph Reasoning: An Uncertainty Aware Perspective

### v1

@@ Abstract
Recently, Knowledge Graphs (KGs) have been successfully coupled with Large Language Models (LLMs) to mitigate their hallucinations and enhance their reasoning capability, e.g., KG-based retrieval-augmented framework. However, current KG-LLM frameworks lack rigorous uncertainty estimation, limiting their reliable deployment in high-stakes applications Directly incorporating uncertainty quantification into KG-LLM frameworks presents challenges due to their complex architectures and the intricate interactions between the knowledge graph and language model components. To address this crucial gap, we propose a new trustworthy KG-LLM framework, UaG (Uncertainty Aware Knowledge-Graph Reasoning), which incorporates uncertainty quantification into the KG-LLM framework. We design an uncertainty-aware multi-step reasoning framework that leverages conformal prediction to provide a theoretical guarantee on the prediction set. To manage the error rate of the multi-step process, we additionally introduce an error rate control module to adjust the error rate within the individual components. Extensive experiments show that UaG can achieve any pre-defined coverage rate while reducing the prediction set/interval size by 40% on average over the baselines.
Introduction
@@
Figure 1: An illustration of uncertainty quantification in the context of knowledge graph question answering. By quantifying a confidence boundary (orange dash), UQ methods can capture more correct answers when facing uncertainties.
@@
The rest of the paper is organized as follows: Section 2 provides the preliminaries; Section 3 introduces our UaG in detail; experiments and ablation studies in Sections 4 and 5; related work in Section 6; lastly, conclusions and future work are presented in Sections 7 and 8.
Preliminiaries
@@
Formally, let $L_{\lambda}:\mathcal{Y}\times\hat{\mathcal{Y}}\rightarrow\mathbb{R}$ be any loss function with a hyperparameter configuration $\lambda\in\Lambda$. Notably, $\lambda$ could be multi-dimensional. Let $\alpha\in\mathbb{R}$ be the user-defined error rate for $L_{\lambda}$ (i.e., $\mathbb{E}(L_{\lambda})\leq\alpha$). Using the calibration set $\mathcal{D}^{\text{cal}}$, LTT then computes a set of valid configurations $\Lambda_{\text{valid}}\in\Lambda$ that satisfies
@@
Substitute $E[L_{\lambda}\mid D_{\text{cal}}]$ in eq. 2, then we have
@@
Experiment
In this section, we conduct extensive experiments on two widely used multi-hop KGQA datasets (Luo et al. 2024; Sun et al. 2023; Lan et al. 2022). Our experiments are designed to rigorously evaluate the uncertainty quantification performance of UaG, employing standard UQ metrics to assess its effectiveness against state-of-the-art baselines. We provide detailed analysis to identify key strength and potential area of improvements. Additional experiments comparing UaG to standard KGQA methods are further detailed in the ablation studies.
Datasets
We evaluate UaG with two widely used benchmark dataset for KGQA: WebQuestionSP (WebQSP) (Yih et al. 2016) and Complex WebQuestions (CWQ) (Talmor and Berant 2018). WebQSP contains up to 2 hop questions, and CWQ contains up to 4 hop questions. Additionally, Freebase (Bollacker et al. 2008) is used as the underlying knowledge graph for both datasets. The dataset statistics are presented in Table 1. Note that for calibration, we use the training partition.
@@
  • APSS (Average Prediction Set Size): This evaluates the effectiveness of the uncertainty quantification model, with smaller average prediction set sizes indicating greater efficiency in selecting the most likely answers.
@@
For the implementation, we set $\delta$ to be 0.05, and use Llama3-8b as our backbone large language model. For the encoder $g$, we use the SentenceTransformer model and pre-train it on our training data. Other implementation details are further documented in the appendix and our code is publicly available [^2] https://github.com/Arstanley/UAG.
@@
Table 2: KGQA Results on WebQSP.
Category Method Hits@1
Fine-tuned TIARA (Shu et al. 2022) 75.2
DeCAF (Yu et al. 2023) 82.1
RoG (Luo et al. 2024) 85.7
Prompting COT (Wei et al. 2022) 39.1
ToG (Sun et al. 2023) 57.4
KD-CoT (Wang et al. 2023b) 63.7
StructGPT (Jiang et al. 2023) 72.6
UaG-Top-1 66.8
UaG-Top-3 73.4
@@
Figure 4: Ablation Studies
@@
As shown in Figure 4(a), pre-training the sentence transformers (specifically for graph traversal) significantly reduces the size of the prediction set. This improvement is likely because fine-tuning at this stage enhances the quality of the initial traversal on the knowledge graph, leading to more accurate retrieval of answers. This process is analogous to better indexing in the context of Retrieval-Augmented Generation (RAG). However, when we pre-trained the large language model backbone, we did not observe a substantial improvement. This outcome aligns with previous findings suggesting that fine-tuning large language models can increase model uncertainty (Ye et al. 2024).
@@
We also experimented with alternative methods for measuring textual similarity, specifically $S_{1}$ as defined in Eq. 4. As shown in Figure 4(b), we found that using cosine similarity significantly improves prediction efficiency while maintaining the user-defined error rate compared to L1 similarity. This improvement can be attributed to cosine similarity’s focus on the angle between vectors, which makes it less sensitive to variations in vector magnitudes. This is particularly helpful in high-dimensional spaces, as it allows cosine similarity to more effectively capture semantic relationships between text representations.
@@
Conclusion
In this paper, we tackle the challenge of uncertainty quantification in knowledge graph question answering by integrating conformal prediction with KG-LLM models. Our architecture leverages the Learn Then Test (LTT) framework for multi-step calibration, delivering reliable results that meet pre-defined error rates while maintaining practical prediction set sizes. Extensive experiments show our method’s effectiveness in balancing accuracy and uncertainty, making it suitable for real-world applications.
@@ Appendix A Appendix
This technical appendix provides additional details and insights to complement the main paper. The rest of the appendix is organized as follows: Section 1 presents the notations used throughout the paper. Section 2 presents the pseudocode for UaG. Section 3 includes the hardware specs and the hyper-parameter settings. Section 4 includes the quantitative experiment results of UaG. Finally, Section 5 includes the Python implementation of UaG.

### v2

@@ Abstract
Recently, Knowledge Graphs (KGs) have been successfully coupled with Large Language Models (LLMs) to mitigate their hallucinations and enhance their reasoning capability, e.g., KG-based retrieval-augmented framework. However, current KG-LLM frameworks lack rigorous uncertainty estimation, limiting their reliable deployment in high-stakes applications Directly incorporating uncertainty quantification into KG-LLM frameworks presents challenges due to their complex architectures and the intricate interactions between the knowledge graph and language model components. To address this crucial gap, we propose a new trustworthy KG-LLM framework, UaG (Uncertainty Aware Knowledge-Graph Reasoning), which incorporates uncertainty quantification into the KG-LLM framework. We design an uncertainty-aware multi-step reasoning framework that leverages conformal prediction to provide a theoretical guarantee on the prediction set. To manage the error rate of the multi-step process, we additionally introduce an error rate control module to adjust the error rate within the individual components. Extensive experiments show that UaG can achieve any pre-defined coverage rate while reducing the prediction set/interval size by 40% on average over the baselines. Our code is available at: https://github.com/Arstanley/UAG
1 Introduction
@@
Figure 1: An illustration of uncertainty quantification in the context of knowledge graph question answering. By quantifying a confidence boundary (dashlines), UQ methods can ensure the accuracy when facing uncertainties.
@@
The rest of the paper is organized as follows: Section 2 provides the preliminaries; Section 3 introduces our UaG in detail; experiments and ablation studies in Sections 4 and 5; related work in Section 6; lastly, conclusions and future work are presented in Sections 7.
2 Preliminiaries
@@
Formally, let $L_{\lambda}:\mathcal{Y}\times\hat{\mathcal{Y}}\rightarrow\mathbb{R}$ be any loss function with a hyperparameter configuration $\lambda\in\Lambda$. Notably, $\lambda$ could be multi-dimensional. Let $\alpha\in\mathbb{R}$ be the user-defined error rate for $L_{\lambda}$ (i.e., $\mathbb{E}(L_{\lambda})\leq\alpha$). Using calibration set $\mathcal{D}^{\text{cal}}$, LTT computes a set of valid configurations $\Lambda_{\text{valid}}\in\Lambda$ satisfying
@@ 3 Method
Substitute $E[L_{\lambda}\mid D_{\text{cal}}]$ in Eq. (2), then we have
@@
4 Experiment
In this section, we conduct extensive experiments on two widely used multi-hop KGQA datasets (Luo et al. 2024; Sun et al. 2023; Lan et al. 2022). Our experiments are designed to rigorously evaluate the uncertainty quantification performance of UaG, employing standard UQ metrics to assess its effectiveness against state-of-the-art baselines.
Datasets
We evaluate UaG with two widely used benchmark dataset for KGQA: WebQuestionSP (WebQSP) (Yih et al. 2016) and Complex WebQuestions (CWQ) (Talmor and Berant 2018). WebQSP contains up to 2 hop questions, and CWQ contains up to 4 hop questions. Additionally, Freebase (Bollacker et al. 2008) is used as the underlying knowledge graph for both datasets. The dataset statistics are shown in Table 1. Note that for calibration, we use the training partition.
@@ 4 Experiment
  • APSS (Average Prediction Set Size):
    This evaluates the effectiveness of the uncertainty quantification model, with smaller average prediction set sizes indicating greater efficiency in selecting the most likely answers.
@@
For the implementation, we set $\delta$ to be 0.05, and use Llama3-8b as our backbone large language model. For the encoder $g$, we use the SentenceTransformer model and pre-train it on our training data. Other implementation details are further documented in the appendix.
@@ 5 Ablation Studies
Table 3: KGQA Results on WebQSP.
Category Method Hits@1
Fine-tuned TIARA (Shu et al. 2022) 75.2
DeCAF (Yu et al. 2023) 82.1
RoG (Luo et al. 2024) 85.7
Prompting COT (Wei et al. 2022) 39.1
ToG (Sun et al. 2023) 57.4
KD-CoT (Wang et al. 2023b) 63.7
StructGPT (Jiang et al. 2023) 72.6
UaG-Top-1 66.8
UaG-Top-3 73.4
@@
Figure 4: Ablation studies where (a) shows the benefits of pre-training the sentence transformer used in graph traversal and (b) shows a comparison when using L1 similarity as opposed to cosine similarity within UaG.
@@
As shown in Figure 4(a), pre-training the sentence transformers (specifically for graph traversal) significantly reduces the size of the prediction set. This improvement is likely because fine-tuning at this stage enhances the quality of the initial traversal on the knowledge graph, leading to more accurate retrieval of answers. This process is analogous to better indexing in the context of Retrieval-Augmented Generation (RAG). This finding aligns with broader trends observed in graph-based question answering systems. However, when we pre-trained the large language model backbone, we did not observe a substantial improvement. This outcome aligns with previous findings suggesting that fine-tuning large language models can increase model uncertainty (Ye et al. 2024).
@@
We also experimented with alternative methods for measuring textual similarity, specifically $S_{1}$ as defined in Eq. (4). As shown in Figure 4(b), we found that using cosine similarity significantly improves prediction efficiency while maintaining the user-defined error rate compared to L1 similarity. This improvement can be attributed to cosine similarity’s focus on the angle between vectors, which makes it less sensitive to variations in vector magnitudes. This is particularly helpful in high-dimensional spaces, as it allows cosine similarity to more effectively capture semantic relationships between text representations.
@@ 6 Related Work
7 Conclusion
In this paper, we tackle the challenge of uncertainty quantification in knowledge graph question answering by integrating conformal prediction with KG-LLM models. Our architecture leverages the Learn Then Test (LTT) framework for multi-step calibration, delivering reliable results that meet pre-defined error rates while maintaining practical prediction set sizes. Extensive experiments show our method’s effectiveness in balancing accuracy and uncertainty, making it suitable for real-world applications, especially those related to high stakes scenarios.
@@ Appendix A Appendix
This technical appendix provides additional details and insights to complement the main paper. The rest of the appendix is organized as follows: Section A1 presents the notations used throughout the paper. Section A2 presents the pseudocode for UaG. Section A3 includes the hardware specs and the hyper-parameter settings. Section A4 includes the quantitative experiment results of UaG. Finally, Section A5 includes the Python implementation of UaG.